An archer fish spies an insect on a stalk of grass

overhanging the pond, and squirts some water at it to

knock it into the water. The insect is 0.30 m

horizontally from the fish and some distance above it.

The fish spits its water at a speed of 3.0 m/s, at an

angle of 75° above horizontal.

(a) How much time does the insect have to react before

the water hits it?

(b) How far above the fish was the insect?

(c) How fast is the water moving when it hits the insect?
Answer #1

(a) t-x/ vo caso 15 t - 0.3 3 coss > = 0.38638 (6) h= Vosmot - I ge? = (3.8m 75°) (0.3863) - 4.9 (0.3863) Th= 0.388 m C) Vx=